# Explainable Reinforcement Learning Agents Using World Models

### v1

@@ Abstract
Explainable AI (XAI) systems have been proposed to help people understand how AI systems produce outputs and behaviors. Explainable Reinforcement Learning (XRL) has an added complexity due to the temporal nature of sequential decision-making. Further, non-AI experts do not necessarily have the ability to alter an agent or its policy. We introduce a technique for using World Models to generate explanations for Model-Based Deep RL agents. World Models predict how the world will change when actions are performed, allowing for the generation of counterfactual trajectories. However, identifying what a user wanted the agent to do is not enough to understand why the agent did something else. We augment Model-Based RL agents with a Reverse World Model, which predicts what the state of the world should have been for the agent to prefer a given counterfactual action. We show that explanations that show users what the world should have been like significantly increase their understanding of the agent’s policy. We hypothesize that our explanations can help users learn how to control the agent’s execution through by manipulating the environment.
@@ 1 Introduction
If an XAI system can help users understand how the agent is responding to the environment, how the policy is making predictions, or how the policy was learned, they may be able to adjust to the agent or the environment, so that the agent’s behavior aligns with our preferences for the policy. Because DRL is applicable to agents and robot planning, many real world applications, from self-driving cars to home personal robotics, will have systems driven by DRL interacting with with users without technical background. Non-AI experts will seek explanations for actions that deviate from their understanding of what the agent should be doing or how it should be doing it Miller (2019). Without an explanation, users can find it difficult to trust the agent’s ability to act safely and reasonably Zelvelder et al. (2021), especially if they have limited artificial intelligence expertise. Even in the case where the agent is operating optimally and without failure, the agent’s optimal behavior may not match the user’s expectations, resulting in confusion and lack of trust. Non-AI experts may also wonder what they can do to influence or change the behavior of the agent if they cannot alter the internal working of the agent or retrain the policy.
@@
Explanation of reinforcement learning systems is especially challenging due to the temporal nature of decision-making—any given decision to perform an action may depend on future expectations as much as, if not more than, current observations. This is made more complicated when accounting for users who are not AI experts and, furthermore, do not have the ability to change the algorithm or retrain the agent. What would an actionable explanation look like for non-AI experts?
@@
In this paper we present a technique for using a World Model (WM) to provide actionable information to non-AI experts of DRL systems. A WM describes the agent’s understanding of the state-transition dynamics of the environment. In Model-Based DRLKaelbling et al. (1996), the agent learns a World Model through trial-and-error interactions with the environment during training. The World Model is used to predict the effects of the agent’s actions, often speeding up policy learning when interactions with the environment are slow or resource intensive. In the scenario in which the agent fails a task or performs an unexpected action, we show that the World Model can be used to generate a counterfactual explanation that shows the user what the agent would have expected the world state to look like in order to have chosen an action preferred by the user. The user can then act upon this information to influence the agent’s actions in the future by affecting change on the environment. This is because the agent’s policy models the relationship between it and the environment with respect to action and the WM makes that relationship explicit. While the user may not be able to change or re-train the policy, the user can alter the agent’s behavior by directly manipulating the environment if the agent has a robust policy.
@@ 4 Preliminaries: Model-Based RL
Reinforcement learning solves sequential decision-making problems. Specifically, it generates a policy, , that maps a state observation to an action such that executing the action optimizes future expected reward if the policy is followed henceforth. Deep reinforcement learning, approximates the policy with a neural network. The policy model is charged with interpreting the state observation with respect to future expected reward. The policy thus captures the dynamics of the agent with respect to the environment and task.
@@ 5 Reverse World Models
The Reverse World Model (RWM) predicts the embedded state . The RWM trains alongside the WM, on a modified copy of the same training data, as follows. First, reward and continuation data is removed. Second, the temporal order of the state-action transition data generated during trials is reversed. DreamerV3 trains its models using an experience replay buffer. DreamerV3 samples chunks of transition data from the replay buffer. When training the RWM only, we simply reverse the order of of the data sampled from the replay buffer. Finally, we shift the actions so that image and the prior action are paired so that when the RWM is trained it predicts the prior embedded state given the future state and prior action.
@@ 6.1 Study Design
We present four scenarios, in each of which the agent fails to complete the fictional coffee recipe for a different reason (see description of scenarios below). For each scenario, the participant must identify what was tampered with in the environment that caused the agent to operate ineffectively. In the Control condition, participants were presented with (a) the video, and (b) four snapshots of states during the agent’s execution. In the Treatment condition, participants were presented with the same information as above, but also (c) each snapshot was paired with a generated image showing what the agent was expecting in order to choose the action that would resulted in correct execution.
@@ 6.1.2 Surveys
The satisfaction survey adapted from Hoffman et al. Hoffman et al. (2023). It consists of seven 5-point Likert scale questions that ask about several key attributes of explanation satisfaction. All questions are framed positively and the answers are supposed to range from ‘Strongly Agree’ to ‘Strongly Disagree’. We mapped these responses as ‘Strongly Agree’=5 down to ‘Strongly Disagree’=1 for simplifying later calculations. We edited the wording of the questions to fit our study while maintaining the same meaning. To have a fair baseline, both the groups were asked these questions about the Error Reports. The error reports presented to the different participant groups differed in that the treatment group received the extra RWM-generated snapshots.
We used the User Trust Survey, also from Hoffman et al. Hoffman et al. (2023). This survey consists of eight 5-point Likert scale survey. All questions are framed positively, except one (“I am wary of the tool”).
Our third survey was is the NASA Task Load Index (TLX) Survey Hart and Staveland (1988), which assesses perceptions of cognitive load. It consists of six questions, each answered on a scale of 21 gradations.
@@ 6.2 Agent Configuration
The agent is our modified DreamerV3 model with reverse world model. It is an Actor-Critic model trained alongside a 25 million parameter FWM and a 25 million parameter RWM. The actor and critic are both MLPs, with learning rates of 3e-5, a batch size of 8 and batch length of 65. The observation input is an image of a 7x7 grid which depicts the agent’s field of view at a given time step. The full training grid for each static environment is 15x15.
The agent never experiences items being moved, obstructed, or removed during training. It is intentionally overfit for purposes of experimental control. While overfit policies are generally unwanted, we required an unnaturally brittle agents that would struggle with the task to construct our experimental conditions; it makes the point that no agent in a complex real-world setting can be perfect.
@@ 7.1 Accuracy
Explanation | No-Explanation | |||
Remove Non-Essential Object | 27.03 | 3.03 | 11.85 | 0.00571* |
Move Essential Object | 86.49 | 33.33 | 12.80 | 0.00001* |
Obstruct Essential Object | 62.16 | 27.27 | 4.38 | 0.00337* |
Remove Essential Object | 83.78 | 42.42 | 7.01 | 0.00034* |
All Scenarios | 64.86 | 26.52 | 5.12 | 0.00001* |
@@ 7.5 Task Completion Time
We recorded the amount of time that participants took to complete each scenario. This encompasses the time they spent reading the status, looking at the video, looking at the images, assembling the response from the pre-populated lists, and writing a short free-text description of how they arrived at the answer. Participants in the Treatment group took on average 3.55 (Std.Dev. 2.57) minutes to complete each scenario. Participants in the Control group took on average 3.68 (Std.Dev. 3.09) minutes to complete each scenario. A t-test indicates that there is no significant difference in completion times between groups (). Table 2 breaks out the time per group and per scenario. Taken along with other results, the explanations in the Treatment Group neither add cognitive load nor add cognitive processing time. Three of four scenarios result in less time, though not significantly so. One scenario took longer on average; this scenario was also significantly harder (per accuracy results in Table 1).
@@
Explanation | No-Explanation | |||
Remove Non-Essential Object | 4.37 (2.23) | 3.80 (2.16) | 1.07 | 0.86 |
Move Essential Object | 3.27 (2.93) | 3.75 (2.57) | -0.72 | 0.24 |
Obstruct Essential Object | 3.65 (2.79) | 4.02 (4.78) | -0.39 | 0.35 |
Remove Essential Object | 2.91 (1.97) | 3.16 (1.88) | -0.53 | 0.29 |
All Scenarios | 3.55 (2.57) | 3.68 (3.09) | -0.38 | 0.35 |
@@ 9 Conclusions
Explainable RL agents for non-AI epxerts presents a significant challenge because the user cannot change the agent or re-train the policy if it doesn’t perform as expected. To explain why an RL agent chose a particular action over another, counterfactual, action, we generate a state in which the agent would have chosen the user’s desired action. To do this, we extend the RL agent with a reverse world model that can predict and generate counterfactual states that preceded the current state, instead of the more common prediction and generation of future counterfactual states.

### v2

@@ Abstract
Explainable AI (XAI) systems have been proposed to help people understand how AI systems produce outputs and behaviors. Explainable Reinforcement Learning (XRL) has an added complexity due to the temporal nature of sequential decision-making. Further, non-AI experts do not necessarily have the ability to alter an agent or its policy. We introduce a technique for using World Models to generate explanations for Model-Based Deep RL agents. World Models predict how the world will change when actions are performed, allowing for the generation of counterfactual trajectories. However, identifying what a user wanted the agent to do is not enough to understand why the agent did something else. We augment Model-Based RL agents with a Reverse World Model, which predicts what the state of the world should have been for the agent to prefer a given counterfactual action. We show that explanations that show users what the world should have been like significantly increase their understanding of the agent’s policy. We hypothesize that our explanations can help users learn how to control the agent’s execution through manipulating the environment.
@@ 1 Introduction
If an XAI system can help users understand how the agent is responding to the environment, how the policy is making predictions, or how the policy was learned, they may be able to adjust to the agent or the environment, so that the agent’s behavior aligns with our preferences for the policy. Because DRL is applicable to agents and robot planning, many real-world applications, from self-driving cars to personal robotics at home, will have systems driven by DRL interacting with users without technical background. Non-AI experts will seek explanations for actions that deviate from their understanding of what the agent should be doing or how it should be doing it Miller (2019). Without an explanation, users can find it difficult to trust the agent’s ability to act safely and reasonably Zelvelder et al. (2021), especially if they have limited artificial intelligence expertise. Even in the case where the agent is operating optimally and without failure, the agent’s optimal behavior may not match the user’s expectations, resulting in confusion and lack of trust. Non-AI experts may also wonder what they can do to influence or change the behavior of the agent if they cannot alter the internal workings of the agent or retrain the policy.
@@
Explanation of reinforcement learning systems is especially challenging due to the temporal nature of decision-making—any given decision to perform an action may depend on future expectations as much as, if not more than, current observations. This becomes more complicated when one accounts for users who are not AI experts and, furthermore, do not have the ability to change the algorithm or retrain the agent. What would an actionable explanation look like for non-AI experts?
@@
In this paper, we present a technique for using a World Model (WM) to provide actionable information to non-AI experts of DRL systems. A WM describes the agent’s understanding of the state-transition dynamics of the environment. In Model-Based DRLKaelbling et al. (1996), the agent learns a World Model through trial-and-error interactions with the environment during training. The World Model is used to predict the effects of the agent’s actions, often speeding up policy learning when interactions with the environment are slow or resource intensive. In the scenario in which the agent fails a task or performs an unexpected action, we show that the World Model can be used to generate a counterfactual explanation that shows the user what the agent would have expected the world state to look like in order to have chosen an action preferred by the user. The user can then act upon this information to influence the agent’s actions in the future by affecting change on the environment. This is because the agent’s policy models the relationship between it and the environment with respect to action, and the WM makes that relationship explicit. While the user may not be able to change or re-train the policy, the user can alter the agent’s behavior by directly manipulating the environment if the agent has a robust policy.
@@ 2 Related Work
Counterfactual explanations, in particular, provide high-level, actionable information which is suitable for non-AI expert users Gajcin and Dusparic (2024b). People naturally explain in “the form ‘Why P rather than Q?’, in which P is the target event and Q is a counterfactual contrast case that did not occur” Miller (2019). Counterfactual (CF) explanations answer the question, what needs to be changed in the input of a computer system such that its output changes to a different one. In sequential decision-making paradigms the output is the choice of a different action than the one the model has taken or wants to take next. In sequential decision-making, counterfactuals can also take the form of hypothetical future action trajectories to show a user what future expected states could be achieved if an alternative action is taken.
Most closely related to our work are CF explanation systems that generate counterfactual states for which the agent would have chosen a different, desired action if those states were encountered Olson et al. (2021); Huber et al. (2023); Samadi et al. (2025). These works employ a second round of training to create a separate, explanation generation model, using action trace data produced by a pre-trained RL agent. This approach has been noted to potentially create counterfactual states that contain the requisite surface features necessary to push the agent to select a new action, but does not guarantee that those states are reachable in the state space Gajcin and Dusparic (2024a). Our technique instead generates explanations from the world model, which is trained as part of the RL agent’s policy training loop. Our technique does not require a separate post-hoc training phase, and the world model—a predictor of successor and predecessor states—is more likely to produce traversable states.
The RACCER system Gajcin and Dusparic (2024a) generates counterfactual states through a search of reachable states for those with the desired properties. It requires access to the execution environment—a simulation, or real-world—to conduct the search. Our technique does not require access to the execution environment, as the world model is a learned surrogate for the environment’s transition dynamics.
COViz Amitai et al. (2024) takes a different approach to counterfactuals and generates states that are predicted to be visited if an alternative action were to be chosen over the actually chosen action. This is potentially useful information to users in evaluating whether an alternative is better than an actual action choice. This approach is complementary to our work, which helps users determine how to induce the agent to take a different action without explicit override.
@@ 4 Preliminaries: Model-Based RL
Reinforcement learning solves sequential decision-making problems. Specifically, it generates a policy, , that maps a state observation to an action such that executing the action optimizes future expected reward if the policy is followed henceforth. Deep reinforcement learning approximates the policy with a neural network. The policy model is charged with interpreting the state observation with respect to future expected reward. The policy thus captures the dynamics of the agent with respect to the environment and task.
@@ 5 Reverse World Models
The Reverse World Model (RWM) predicts the embedded state . The RWM trains alongside the WM, on a modified copy of the same training data, as follows. First, reward and continuation data is removed. Second, the temporal order of the state-action transition data generated during trials is reversed. DreamerV3 trains its models using an experience replay buffer. DreamerV3 samples chunks of transition data from the replay buffer. When training the RWM only, we simply reverse the order of the data sampled from the replay buffer. Finally, we shift the actions so that image and the prior action are paired so that when the RWM is trained it predicts the prior embedded state given the future state and prior action.
@@ 6.1 Study Design
We present four scenarios, in each of which the agent fails to complete the fictional coffee recipe for a different reason (see description of scenarios below). For each scenario, the participant must identify what was tampered with in the environment that caused the agent to operate ineffectively. In the Control condition, participants were presented with (a) the video, and (b) four snapshots of states during the agent’s execution. In the Treatment condition, participants were presented with the same information as above, but also (c) each snapshot was paired with a generated image showing what the agent was expecting in order to choose the action that would have resulted in correct execution.
@@ 6.1.2 Surveys
The satisfaction survey is adapted from Hoffman et al. Hoffman et al. (2023). It consists of seven 5-point Likert scale questions that ask about several key attributes of explanation satisfaction. All questions are framed positively, and the answers are supposed to range from ‘Strongly Agree’ to ‘Strongly Disagree’. We mapped these responses as ‘Strongly Agree’=5 down to ‘Strongly Disagree’=1 for simplifying later calculations. We edited the wording of the questions to fit our study while maintaining the same meaning. To have a fair baseline, both the groups were asked these questions about the Error Reports. The error reports presented to the different participant groups differed in that the treatment group received the extra RWM-generated snapshots.
We used the User Trust Survey, also from Hoffman et al. Hoffman et al. (2023). This survey consists of eight 5-point Likert scale questions. All questions are framed positively, except one (“I am wary of the tool”).
Our third survey is the NASA Task Load Index (TLX) Survey Hart and Staveland (1988), which assesses perceptions of cognitive load. It consists of six questions, each answered on a scale of 21 gradations.
@@ 6.2 Agent Configuration
The agent is our modified DreamerV3 model with a reverse world model. It is an Actor-Critic model trained alongside a 25 million parameter FWM and a 25 million parameter RWM. The actor and critic are both MLPs, with learning rates of 3e-5, a batch size of 8 and a batch length of 65. The observation input is an image of a 7x7 grid which depicts the agent’s field of view at a given time step. The full training grid for each static environment is 15x15.
The agent never experiences items being moved, obstructed, or removed during training. It is intentionally overfit for purposes of experimental control. While overfit policies are generally unwanted, we required an unnaturally brittle agent that would struggle with the task to construct our experimental conditions; it makes the point that no agent in a complex real-world setting can be perfect.
@@ 7.1 Accuracy
| Control | Treatment | |||
| Remove Non-Essential Object | 3.03 | 27.03 | 11.85 | 0.00571* |
| Move Essential Object | 33.33 | 86.49 | 12.80 | 0.00001* |
| Obstruct Essential Object | 27.27 | 62.16 | 4.38 | 0.00337* |
| Remove Essential Object | 42.42 | 83.78 | 7.01 | 0.00034* |
| All Scenarios | 26.52 | 64.86 | 5.12 | 0.00001* |
@@ 7.5 Task Completion Time
We recorded the amount of time that participants took to complete each scenario. This encompasses the time they spent reading the status, looking at the video, looking at the images, assembling the response from the pre-populated lists, and writing a short free-text description of how they arrived at the answer. Participants in the Treatment group took on average 3.55 (Std.Dev. 2.57) minutes to complete each scenario. Participants in the Control group took on average 3.68 (Std.Dev. 3.09) minutes to complete each scenario.
A t-test indicates that there is no significant difference in completion times between groups (). Table 2 breaks out the time per group and per scenario. Taken along with other results, the explanations in the Treatment Group neither add cognitive load nor add cognitive processing time. Three of four scenarios result in less time, though not significantly so. One scenario took longer on average; this scenario was also significantly harder (per accuracy results in Table 1).
@@
| Control | Treatment | |||
| Remove Non-Essential Object | 3.80 (2.16) | 4.37 (2.23) | 1.07 | 0.86 |
| Move Essential Object | 3.75 (2.57) | 3.27 (2.93) | -0.72 | 0.24 |
| Obstruct Essential Object | 4.02 (4.78) | 3.65 (2.79) | -0.39 | 0.35 |
| Remove Essential Object | 3.16 (1.88) | 2.91 (1.97) | -0.53 | 0.29 |
| All Scenarios | 3.68 (3.09) | 3.55 (2.57) | -0.38 | 0.35 |
@@ 9 Conclusions
Explainable RL agents for non-AI experts presents a significant challenge because the user cannot change the agent or re-train the policy if it doesn’t perform as expected. To explain why an RL agent chose a particular action over another, counterfactual, action, we generate a state in which the agent would have chosen the user’s desired action. To do this, we extend the RL agent with a reverse world model that can predict and generate counterfactual states that preceded the current state, instead of the more common prediction and generation of future counterfactual states.